OpenStudy (anonymous):

Find the quotient. 6 /2 /7÷ 2/3/4

OpenStudy (ranga):

Dividing by a fraction is same as multiplying by its reciprocal: a / b / c = a * c / b

OpenStudy (anonymous):

I tried

OpenStudy (ranga):

6 / 2 / 7 = 6 * 7/2 = 3 * 7 = 21 2 / 3 / 4 = 2 * 4/3 = 8/3 6 /2 /7÷ 2/3/4 = 21 / 8/3 = 21 * 3/8 = 63/8 = 7 & 7/8.